OpenStudy (anonymous):

Solve the equation -5y -9 = -(y - 1)

OpenStudy (anonymous):

-5y -9 = 1 - y

OpenStudy (anonymous):

y = -5/2

OpenStudy (anonymous):

I have -2/5

OpenStudy (anonymous):

yea thats it

OpenStudy (anonymous):

Thank you @sammaraq60